Old Country Links, Inc., produces sausages in three production departments—Mixing, Casing and Curing, and Packaging. In the Mixing Department, meats are prepared and ground and then mixed with spices. The spiced meat mixture is then transferred to the Casing and Curing Department, where the mixture is force-fed into casings and then hung and cured in climate-controlled smoking chambers. In the Packaging Department, the cured sausages are sorted, packed, and labeled. The company uses the weighted-average method in its process costing system. Data for September for the Casing and Curing Department follow:

Percent Completed
Units Mixing Materials Conversion
Work in process inventory, September 1 5 100 % 60 % 50 %
Work in process inventory, September 30 5 100 % 20 % 10 %


Mixing Materials Conversion
Work in process inventory, September 1 $ 13,240 $ 25 $ 4,005
Cost added during September $ 291,680 $ 21,035 $ 202,200


Mixing cost represents the costs of the spiced meat mixture transferred in from the Mixing Department. The spiced meat mixture is processed in the Casing and Curing Department in batches; each unit in the above table is a batch and one batch of spiced meat mixture produces a set amount of sausages that are passed on to the Packaging Department. During September, 116 batches (i.e., units) were completed and transferred to the Packaging Department.

Required:

1. Determine the Casing and Curing Department's equivalent units of production for mixing, materials, and conversion for the month of September.

2. Compute the Casing and Curing Department's cost per equivalent unit for mixing, materials, and conversion for the month of September.

3. Compute the Casing and Curing Department's cost of ending work in process inventory for mixing, materials, conversion, and in total for September.

4. Compute the Casing and Curing Department's cost of units transferred out to the Packaging Department for mixing, materials, conversion, and in total for September.

5. Prepare a cost reconciliation report for the Casing and Curing Department for September.

Determine the Casing and Curing Department's equivalent units of production for mixing, materials, and conversion for the month of September. (Round your intermediate calculations and final answers to 1 decimal place.)

Mixing Materials Conversion
Equivalent units of production

Mixing Materials Conversion
Cost per equivalent unit

Compute the Casing and Curing Department's cost of ending work in process inventory for mixing, materials, conversion, and in total for September. (Round your intermediate calculations to 1 decimal place and your final answers to the nearest whole dollar amount.)

Mixing Materials Conversion Total
Cost of ending work in process inventory

Compute the Casing and Curing Department's cost of units transferred out to the Packaging Department for mixing, materials, conversion, and in total for September. (Round your intermediate calculations to 1 decimal place and your final answers to the nearest whole dollar amount.)

Mixing Materials Conversion Total
Cost of units transferred out

Prepare a cost reconciliation report for the Casing and Curing Department for September. (Round your intermediate calculations to 1 decimal place and your final answers to the nearest whole dollar amount.)

Casing and Curing Department
Cost Reconciliation
Costs to be accounted for:
Total cost to be accounted for $0
Costs accounted for as follows:
Total cost accounted for $0
Answer #1

1. Casing and Curing Department's equivalent units of production are:

Formula = Units Completed and transferred to next department + (Closing Work-in-Process Inventory * Percentage complete of Closing Stock)

Mixing = 116 + 5 = 121 units

Material = 116 + 20%*5 = 116 + 1 = 117 units

Conversion = 116 + 10%*5 = 116 + 0.5 = 116.50 units

2. Casing and Curing Department's cost per equivalent unit are:

Formula = (Cost of Beginning Work in Process Inventory + Cost added during the year)/Equivalent Units during the period

Mixing = $(13,240 + 291,680)/ 121 = $2,520

Material = $(25 + 21,035)/ 117 = $180

Conversion = $(4,005 + 202,200)/ 116.50 = $1,770

3. Casing and Curing Department's cost of ending work in process inventory are:

Formula = Cost per unit (as calculated in point 2) * Closing Work-in-process inventory

Mixing = $2,520 * 5 = $12,600

Material = $180 * 1 = $180

Conversion = $1,770 * 0.5 = $885

Total Cost = $(12,600 + 180 + 885) = $13,665

4. Casing and Curing Department's cost of units transferred out to the Packaging Department are:

Formula = Cost per unit (as calculated in point 2) * 116 (units transferred to the Packaging Department)

Mixing = $2,520 * 116 = $292,320

Material = $180 * 116 = $20,880

Conversion = $1,770 * 116 = $205,320

Total Cost = $(292,320 + 20,880 + 205,320) = $518,520

5. Cost Reconciliation Report for the Casing and Curing Department for September:

Cost to be accounted for:

Cost of beginning work-in-process inventory

$ 17,270

Costs added during the period

$ 514,915

Total cost accounted for September

$532,185

Cost accounted for as follows:

Cost of ending work-in-process inventory

$13,665

Cost of units transferred to Packaging Department

$518,520

Total cost accounted for September

$532,185